Find the linearization of f(x)=1/x^2 at x=2 and use it to approximate the value of 1/1.98^2
Answer #1

Linearization of a function f (x at x=a is defined by

L ()f(a)f (a) (x -6

---------------------------------------------------------

Here

f (x) -2 = x

f (x) dr 2 ()

--2a3

At x=2

f (2) 22

f (2)-2 x 2-3 1 4

Linearization is given by

L ()f(2)f (2) (-2)

=\frac{1}{4}-\frac{1}{4}\left ( x-2 \right )

Therefore

To find value of \frac{1}{1.98^{2}} substitute x=1.98

f\left ( x \right )\approx L\left ( x \right )=\frac{1}{4}-\frac{1}{4}\left ( x-2 \right )

f\left ( 1.98 \right )\approx L\left ( 1.98 \right )=\frac{1}{4}-\frac{1}{4}\left ( 1.98-2 \right )=0.255

That is

.982 0.255

-------------------------------------------------------------